# Explainable Light-Weight Deep Learning Pipeline for Improved Drought Stress Identification in Potato

### v1

@@ Abstract
Early identification of drought stress in crops is vital for implementing effective mitigation measures and reducing yield loss. Non-invasive imaging techniques hold immense potential by capturing subtle physiological changes in plants under water deficit. Sensor based imaging data serves as a rich source of information for machine learning and deep learning algorithms, facilitating further analysis aimed at identifying drought stress. While these approaches yield favorable results, real-time field applications requires algorithms specifically designed for the complexities of natural agricultural conditions. Our work proposes a novel deep learning framework for classifying drought stress in potato crops captured by UAVs in natural settings. The novelty lies in the synergistic combination of a pre-trained network with carefully designed custom layers. This architecture leverages the pre-trained network’s feature extraction capabilities while the custom layers enable targeted dimensionality reduction and enhanced regularization, ultimately leading to improved performance. A key innovation of our work involves the integration of Gradient-Class Activation Mapping (Grad-CAM), an explainability technique. Grad-CAM sheds light on the internal workings of the deep learning model, typically referred to as a ”black box.” By visualizing the model’s focus areas within the images, Grad-CAM fosters interpretability and builds trust in the model’s decision-making process. Our proposed framework achieves superior performance, particularly with the DenseNet121 pre-trained network, reaching a precision of 98% to identify the stressed class with an overall accuracy of 90%. Comparative analysis of existing state-of-the-art object detection algorithms reveals the superiority of our approach in significantly higher precision and accuracy. Thus, our explainable deep learning framework offers powerful approach for drought stress identification with high accuracy and actionable insights.
@@ 2.2 Model based on Transfer Learning
2.3 Gradient-weighted Class Activation Mapping
Convolutional layers naturally preserve spatial details that are typically lost in fully-connected layers. Therefore, it’s reasonable to anticipate that the final convolutional layers offer the most optimal balance between capturing high-level semantics and retaining intricate spatial information. Neurons within these layers specialize in identifying class-specific features within an image, such as distinct object parts. Grad-CAM [29] utilizes gradient information directed into the last convolutional layer of the CNN to determine the importance of each neuron for a specific decision by visualizing the area of interest related to the decision. Thus, It provides insights into the decision-making process of the model by highlighting the regions that contribute most to its predictions.
The mathematical background of Grad-CAM involves the following concepts:
  • Gradient Ascent: Gradient ascent updates the input image iteratively to maximize the model’s output score for a specific class. Mathematically, it can be expressed as:
    $\text{Input}_{\text{new}}=\text{Input}_{\text{old}}+\alpha\cdot\nabla_{\text{Input}}\text{Output}$
@@ 2.3 Gradient-weighted Class Activation Mapping
    • $\text{Input}_{\text{new}}$ is the updated input image.
    • $\text{Input}_{\text{old}}$ is the previous input image.
    • $\alpha$ is the step size (learning rate).
    • $\nabla_{\text{Input}}\text{Output}$ is the gradient of the model’s output with respect to the input image.
  • Recording Operations: During the forward pass of the model, all operations applied to the input image are recorded. This can be represented as follows:
    Forward $\text{Forward Pass}:\text{Operations}=f(\text{Input})$
    where $f(\cdot)$ represents the operations performed by the model.
  • Computing Gradients: After the operations are recorded, gradients of the model’s output with respect to the input image are computed using the chain rule of calculus:
    $\nabla_{\text{Input}}\text{Output}=\frac{\partial\text{Output}}{\partial\text{Input}}$
    This gradient indicates how a small change in each pixel of the input image would affect the model’s output.
The Grad-CAM process involves the following steps:
  • Model Loading and Modification: The trained model, including its weights and all parameters, is loaded. Subsequently, the model undergoes modification by removing the fully connected layers to access the necessary feature maps.
  • Gradient Computation: Using TensorFlow’s GradientTape, gradients of the model’s output with respect to the input image are computed. This step involves performing a forward pass through the model to obtain the output and then computing gradients using back propagation.
  • Heat map Generation: The computed gradients are used to generate a heat map that highlights the important regions of the input image. The heat map represents the sensitivity of the model’s output to changes in different parts of the image.
@@ 3.2 Performance of the Model
We investigated four pre-trained networks individually as part of the proposed deep learning pipeline. While EfficientNetB0 and NASNetMobile achieved high training accuracies of 99.38% and 98.47%, respectively, their validation and test accuracies were notably lower, indicating potential weaknesses as evidenced by their loss and accuracy learning curves which is discussed later in the section. In contrast, MobileNet demonstrated impressive performance with a training accuracy of 99.81%, validation accuracy of 99.33%, and a test accuracy of 88.72%, coupled with a low validation loss of 0.033. Similarly, DenseNet121 showcased robust performance across training, validation, and test sets, achieving a trainng accuracy of 99.56%, a validation accuracy of 99.01% and a test accuracy of 89.86%. Overall, DenseNet121 emerged as the best-performing model among those investigated, boasting the highest test accuracy, closely followed by MobileNet. The comparative performance is summarized in Table 1. Epochs in training are chosen based on observing the convergence pattern of the model, typically by monitoring performance metrics on a validation dataset. The training process continues until the model’s performance on the validation set plateaus or starts to degrade, indicating convergence and preventing over-fitting. The deep learning pipeline was trained with EfficientNetB0, MobileNet, DenseNet121, and NASNetMobile for 30, 60, 50, and 30 epochs, respectively. The optimal performance for each model was achieved at epochs 30, 59, 49, and 28, correspondingly.
@@
The trained models underwent evaluation using a distinct test set comprising 1135 images, not utilized during the model training phase. The confusion matrices, depicted in Fig. 8, were generated by the model employing various pre-trained networks. The values within each confusion matrix were arranged according to the layout shown in Fig. 4. Following the similar pattern observed in the previous learning curves, the EfficientNetB0 and NASNetMobile showed the poorest performance on the test dataset. For EfficientNetB0, analysis of the confusion matrix (Fig. 8(a)) reveals a total of 295 misclassified instances out of 1135 predictions, comprised of 138 false positives (FP) and 157 false negatives (FN). This results in a mis-classification rate of 26%. In contrast, the confusion matrix for NASNetMobile (Fig. 8(d)) indicates an anomalous behavior where the model correctly identifies all stressed images but fails to recognize any healthy ones. In the case of EfficientNetB0, the higher misclassification rate suggests suboptimal performance across both classes. Conversely, NASNetMobile’s performance is characterized by a notable bias towards the ”stressed” class, resulting in a complete oversight of the ”healthy” class. Both of these scenarios are deemed undesirable, rendering the models ineffective for their intended purpose. On the other hand, both MobileNet and DenseNet121 achieve very low mis-classification rates between healthy and stressed classes, as shown by the minimal Type I and Type II errors in their respective confusion matrices (Fig. 8(b) and Fig. 8(c)). This translates to high overall accuracies of 88.72% for MobileNet and 89.86% for DenseNet121. DenseNet121 emerges as the superior model conceptually based on the learning curves and confusion matrix behavior. It generalizes better on unseen data and exhibits a clearer distinction between the healthy and stressed classes.
@@ 3.4 Comparison of the models
We are evaluating the performance metrics of both state-of-the-art object detection algorithms and our newly proposed classifier as they are tasked with handling identical two classes, utilizing the same dataset and bounding boxes. The higher the values in the precision and recall columns, the better the model’s performance in classifying that particular class (stressed or healthy). Table 2 presents the performance metrics of various models as assessed by [25], which are then compared with the performance of the proposed model. The proposed pipeline with DenseNet121 demonstrates the significantly higher precision for both stressed and healthy instances among the evaluated models, achieving values of 0.976 and 0.795, respectively. While it also boasts the highest recall (i.e. 0.960) for the stressed class compared to all other models, its recall for the stressed class is slightly lower than that of Yolo v3. Yolo v3, while having the highest Recall for stressed plants (0.882), has a lower Precision (0.407) compared to other models, indicating it might be classifying some healthy plants as stressed. This indicates that the proposed method generally outperforms these models in accurately identifying stressed and healthy conditions of crops. This observation is further supported by the histogram shown in Fig. 10.
@@
Proposed Pipeline(with DenseNet121) 0.976 0.865 0.795 0.960

### v3

@@ Abstract
Early identification of drought stress in crops is vital for implementing effective mitigation measures and reducing yield loss. Non-invasive imaging techniques hold immense potential by capturing subtle physiological changes in plants under water deficit. Sensor based imaging data serves as a rich source of information for machine learning and deep learning algorithms, facilitating further analysis aimed at identifying drought stress. While these approaches yield favorable results, real-time field applications requires algorithms specifically designed for the complexities of natural agricultural conditions. Our work proposes a novel deep learning framework for classifying drought stress in potato crops captured by UAVs in natural settings. The novelty lies in the synergistic combination of a pre-trained network with carefully designed custom layers. This architecture leverages the pre-trained network’s feature extraction capabilities while the custom layers enable targeted dimensionality reduction and enhanced regularization, ultimately leading to improved performance. A key innovation of our work involves the integration of Gradient-Class Activation Mapping (Grad-CAM), an explainability technique. Grad-CAM sheds light on the internal workings of the deep learning model, typically referred to as a ”black box.” By visualizing the model’s focus areas within the images, Grad-CAM fosters interpretability and builds trust in the model’s decision-making process. Our proposed framework achieves superior performance, particularly with the DenseNet121 pre-trained network, reaching a precision of 97% to identify the stressed class with an overall accuracy of 91%. Comparative analysis of existing state-of-the-art object detection algorithms reveals the superiority of our approach in significantly higher precision and accuracy. Thus, our explainable deep learning framework offers powerful approach for drought stress identification with high accuracy and actionable insights.
@@ 2.2 Model based on Transfer Learning
2.3 Explaining the Model
Convolutional layers naturally preserve spatial details that are typically lost in fully-connected layers. Therefore, it’s reasonable to anticipate that the final convolutional layers offer the most optimal balance between capturing high-level semantics and retaining intricate spatial information. Neurons within these layers specialize in identifying class-specific features within an image, such as distinct object parts. Grad-CAM [29] utilizes gradient information directed into the last convolutional layer of the CNN to determine the importance of each neuron for a specific decision by visualizing the area of interest related to the decision. Thus, It provides insights into the decision-making process of the model by highlighting the regions that contribute most to its predictions. We are proposing a strategy, based on Grad-CAM to improve model transparency by pinpointing the factors that trigger stress responses.
The following steps are involved in the suggested explainable approach, which takes its cue from GRAD-CAM:
  1. 1.
    Forward Pass: The model output $\theta$ is computed by performing a forward pass through the deep learning model, represented as:
    $\theta=\phi(\xi)$
@@ 2.3 Explaining the Model
    • $\theta$ represents the model output.
    • $\phi(\cdot)$ represents the deep learning model.
    • $\xi$ represents the input image.
  2. 2.
    Compute Gradients: The gradients of the model output with respect to the input image are calculated using TensorFlow’s GradientTape, represented as:
    $\nabla_{\xi}\theta=\frac{\partial\theta}{\partial\xi}$
    where:
    • $\nabla_{\xi}\theta$ represents the gradients of the model output with respect to the input image.
    • $\frac{\partial\theta}{\partial\xi}$ represents the partial derivatives of the model output with respect to the input image.
  3. 3.
    Gradient Visualization: The absolute gradients are computed and visualized as a heatmap, represented as:
    $\text{Heatmap}=\text{abs}(\nabla_{\xi}\theta)$
    where:
    • Heatmap represents the heatmap visualization of the gradients.
[…]
  4. 4.
    Standardization: The heatmap is optionally standardized by subtracting the mean and dividing by the standard deviation to improve visualization, represented as:
    $\text{Heatmap}_{\text{std}}=\frac{\text{Heatmap}-\mu}{\sigma}$
    where:
    • $\text{Heatmap}_{\text{std}}$ represents the standardized heatmap.
    • $\mu$ represents the mean of the heatmap.
    • $\sigma$ represents the standard deviation of the heatmap.
  5. 5.
    Plotting: Finally, the input image and the heatmap are plotted side by side for visualization.
Thus, the explainable approach based on Grad-CAM leverages the strength by analyzing gradients to pinpoint image regions crucial for the model’s decisions, offering valuable insights into what triggers the model’s stress responses.
@@ 3.2 Performance of the Model
We investigated four pre-trained networks individually as part of the proposed deep learning pipeline. While EfficientNetB0 and NASNetMobile achieved high training accuracies of 99.38% and 98.47%, respectively, their validation and test accuracies were notably lower, indicating potential weaknesses as evidenced by their loss and accuracy learning curves which is discussed later in the section. In contrast, MobileNet demonstrated impressive performance with a training accuracy of 99.81%, validation accuracy of 99.33%, and a test accuracy of 88.72%, coupled with a low validation loss of 0.033. Similarly, DenseNet121 showcased robust performance across training, validation, and test sets, achieving a training accuracy of 99.69%, a validation accuracy of 98.86% and a test accuracy of 90.75%. Overall, DenseNet121 emerged as the best-performing model among those investigated, boasting the highest test accuracy, closely followed by MobileNet. The comparative performance is summarized in Table 1. Epochs in training are chosen based on observing the convergence pattern of the model, typically by monitoring performance metrics on a validation dataset. The training process continues until the model’s performance on the validation set plateaus or starts to degrade, indicating convergence and preventing over-fitting. The deep learning pipeline was trained with EfficientNetB0, MobileNet, DenseNet121, and NASNetMobile for 30, 60, 60, and 30 epochs, respectively. The optimal performance for each model was achieved at epochs 30, 59, 55, and 28, correspondingly.
@@
The trained models underwent evaluation using a distinct test set comprising 1135 images, not utilized during the model training phase. The confusion matrices, depicted in Fig. 8, were generated by the model employing various pre-trained networks. The values within each confusion matrix were arranged according to the layout shown in Fig. 4. Following the similar pattern observed in the previous learning curves, the EfficientNetB0 and NASNetMobile showed the poorest performance on the test dataset. For EfficientNetB0, analysis of the confusion matrix (Fig. 8(a)) reveals a total of 295 misclassified instances out of 1135 predictions, comprised of 138 false positives (FP) and 157 false negatives (FN). This results in a mis-classification rate of 26%. In contrast, the confusion matrix for NASNetMobile (Fig. 8(d)) indicates an anomalous behavior where the model correctly identifies all stressed images but fails to recognize any healthy ones. In the case of EfficientNetB0, the higher misclassification rate suggests suboptimal performance across both classes. Conversely, NASNetMobile’s performance is characterized by a notable bias towards the ”stressed” class, resulting in a complete oversight of the ”healthy” class. Both of these scenarios are deemed undesirable, rendering the models ineffective for their intended purpose. On the other hand, both MobileNet and DenseNet121 achieve very low mis-classification rates between healthy and stressed classes, as shown by the minimal Type I and Type II errors in their respective confusion matrices (Fig. 8(b) and Fig. 8(c)). This translates to high overall accuracies of 88.72% for MobileNet and 90.75% for DenseNet121. DenseNet121 emerges as the superior model conceptually based on the learning curves and confusion matrix behavior. It generalizes better on unseen data and exhibits a clearer distinction between the healthy and stressed classes.
@@ 3.4 Comparison of the models
We are assessing the performance metrics of both state-of-the-art object detection algorithms and our newly proposed classifier with explainability features. Both systems are tasked with identifying and distinguishing between the same two classes, utilizing identical datasets and bounding boxes. This comparison is particularly relevant because the localization aspect inherent in object detection algorithms aligns with our proposed approach, which involves pinpointing areas of stress. The evaluation is centered on precision and recall metrics to ascertain their performance and effectiveness in accurately detecting instances of the target classes. Greater precision and recall values indicate superior model performance in classifying specific classes, whether stressed or healthy. Table 2 presents the performance metrics of various models as assessed by [25], which are then compared with the performance of the proposed model. The proposed pipeline with DenseNet121 demonstrates the significantly higher precision for both stressed and healthy instances among the evaluated models, achieving values of 0.967 and 0.820, respectively. While it also boasts the highest recall (i.e. 0.887) for the stressed class compared to all other models, its recall for the stressed class is slightly higher than that of Yolo v3. Yolo v3, while having the impressive recall for stressed plants (0.882), has a lower precision (0.407) compared to other models, indicating it might be classifying some healthy plants as stressed. This indicates that the proposed method generally outperforms these models in accurately identifying stressed and healthy conditions of crops. This observation is further supported by the histogram shown in Fig. 10.
@@
Proposed Pipeline(with DenseNet121) 0.967 0.887 0.820 0.945
@@ Acknowledgements
Ethical Statement
We like to declare that all the authors mentioned in the manuscript have agreed for authorship, read and approved the manuscript, and given consent for submission and subsequent publication of the manuscript. The communicating author will be responsible for the integrity of the manuscript (including ethics, data handling, reporting of results, and study conduct),  would communicate with the journal if any technical clarifications related to the manuscript are required, and would handle similar responsibilities. We confirm that this work is original and has not been published elsewhere, nor is it currently under consideration for publication elsewhere We also declare that the study conducted do not involve human subjects and or animals and therefore, do not require any institutional ethical approval. Authors also declare that there is no any financial and non-financial competing interests with anyone concerned.